OpenStudy (anonymous):

Given f(x) = the quantity of 5x plus 1, divided by 2, solve for f-1(8).

OpenStudy (anonymous):

3 7 10 20

OpenStudy (anonymous):

@iGreen

OpenStudy (harsha19111999):

Solve for f-1(8) or f=8?

OpenStudy (anonymous):

idk how to?

OpenStudy (anonymous):

@iGreen can u help me

OpenStudy (harsha19111999):

I'm asking the last part of the question.I didn't understand it. If x=8, then you get 20

OpenStudy (igreen):

Is this it? \(f(x) = \Large\frac{5x + 1}{2}\)

OpenStudy (anonymous):

yeah

OpenStudy (igreen):

\(8 = \Large\frac{5x + 1}{2}\) Is this it? @Harsha19111999

OpenStudy (anonymous):

I think its 20 I have two more questions

OpenStudy (harsha19111999):

Yeah it is but the last part is it f(x)=8 or f-1(8)?

OpenStudy (harsha19111999):

If its is f(x)=8 the answer will be 3

OpenStudy (igreen):

Yeah